1.How to draw/interpret the Napier’s circle?
2.What is the meaning of Napier’s Formula?
3. How to formulate Napier’s Formula?
4.How do you derive formula from the formulated Napier’s Formula?
1
Expert's answer
2020-10-08T15:57:45-0400

1.Draw a right triangle on a sphere and label the sides "a, b" and "c"  where "c"  is the hypotenuse. Let A be the angle opposite side "a" , B the angle opposite side "b" , and C the right angle opposite the hypotenuse "c" .


2.Then Napier’s Formula has two rules: The sine of a part is equal to the product of the tangents of the two adjacent parts. The sine of a part is equal to the product of the cosines of the two opposite parts.


3.In any triangle ABC,


(i) "tan (\frac{B−C} {2} ) = (\frac{b−C} {b+c} ) cot\frac{A} {2}"


(ii) "tan (\frac{C−A} {2} ) = (\frac{c−a} {c+a} ) cot \frac{B} {2}"


(iii) "tan (\frac{A−B} {2} ) = (\frac{a−b} {a+b} ) cot \frac{C} {2}"


4.Considering "c" , whose non-adjacent parts are 𝑎

and 𝑏.By rule (b), cos c = sin 𝑎sin 𝑏, i.e.,cos c = cos a cos b.This becomes obvious from the cosine rule (1), since γ = 90° and, thus, sin a sin b cos γ = 0. In the same way rule (a) for the adjacent parts α and β of c reads

cos c = cot α cot β.